# FedNoisy: Federated Noisy Label Learning Benchmark

### v1

@@ Appendix B Data Partitions
Given a data sample $(x,y)$, non-IID indicates the joint probability distribution $\mathcal{P}(x,y)$ is different for different client $i$ and $j$ ($i,j\in[K]$, $i\neq j$), that is, $\mathcal{P}_{i}(x,y)\neq\mathcal{P}_{j}(x,y)$. By Bayes’ theorem, $\mathcal{P}(x,y)$ can be rewritten as $\mathcal{P}(y|x)\mathcal{P}(x)$ and $\mathcal{P}(x|y)\mathcal{P}(y)$. Based on previous research [50, 56, 46], we present IID and three representative non-IID schemes in FedNoisy. In this section, we first introduce IID partition, then partition methodologies for non-IID cases.
@@
IID   In IID, data distributions on each client are independent and identical. Following [50], given a global dataset with $N$ training samples, we first randomly shuffle it, then partition it over $K$ clients, each receiving $\lfloor\frac{N}{K}\rfloor$ samples. Note that this partition keeps the original class ratio in the global dataset. If a class imbalance problem exists in the global dataset, then each client has the same class imbalance issue on its local dataset. We use iid to denote the IID partition.
@@
Quantity skew   In quantity distribution skew, different clients can hold different amounts of data, while keeping the same class distribution. For example, large academic hospitals usually have larger datasets compared with small community hospitals [66]. Dirichlet distribution [46] and Log-Normal distribution [67] can be used to simulate this Non-IID partition. Here we use Dirichlet distribution following [46]. Given $K$ clients, we draw $K$-dim vector $\mathbf{q}\sim\mathrm{Dir}_{K}(\alpha)$ and allocate $\lfloor q_{k}N\rfloor$ samples from global dataset $\mathcal{D}$ to client $k$ for $k=1,\ldots,K$. The parameter $\alpha>0$ can regulate the imbalance level of the quantity skew: a larger $\alpha$ will even lead to a partition result close to the IID partition. We use noniid-quantity=$\alpha$ to denote quantity distribution skew with parameter value $\alpha$.
@@
Label distribution skew   In label distribution skew, the prior distributions $\mathcal{P}(y)$ may vary across clients, while $\mathcal{P}(x|y)$ is shared in the federated system. For example, primary angle closure glaucoma (PACG) is much more common than primary open-angle glaucoma (POAG) in the Asian population than African, Hispanic population [68]. In FedNoisy, we include two common practices for label distribution skew, namely distribution-based label distribution skew and quantity-based label distribution skew.
@@
  • Dirichlet label skew uses Dirichlet distribution to determine the class distribution on clients [69, 70, 46]. Given $K$ clients and $C$ classes, we draw a $K$-dim vector $\mathbf{q}_{i}$ from $\mathrm{Dir}_{K}(\alpha)$ for each class $i\in[C]$ and allocate $q_{ik}$ proportion of global dataset with label $i$ to local client $k$. In addition to various overall sample sizes, the clients also have varying class distributions. A larger $\alpha$ encourages the class distribution among clients close to a uniform distribution. We use noniid-labeldir=$\alpha$ to denote distribution-based label distribution skew with parameter $\alpha$.
@@
  • Label-quantity skew assigns samples of specified number of classes on each client [50, 46, 71]. Here we follow the strategy in [46]. Given the number of classes on each client as $c$, we first assign $c$ different classes to each client and then equally divide samples of each class according to the number of clients assigned to them. When $c$ equals the total class number in the federation, the partition is equivalent to the IID partition. We use noniid-#label=$c$ to denote quantity-based label distribution skew with parameter $c$.
@@ Appendix C Fair comparison between globalized and localized noise scene
Figure 6: Overall noisy label ratio in localized symmetric noise $\mathcal{U}(\varepsilon_{local}-\sigma,\varepsilon_{local}+\sigma)$ on CIFAR-10 with 10 clients. From left to right: iid, noniid-labeldir, noniid-#label and noniid-labeldir partition.
@@ Appendix D Gradient norm under localized noise setting
Figure 7: Gradient norm of global model for clean and localized noise $\varepsilon_{k}\sim\mathcal{U}(0.3,0.5)$.
@@ Appendix E Results on other datasets
E.1 MNIST results on FedAvg
@@
Table 4: Test accuracy of MNIST on 10 clients using FedAvg with CNN
@@ E.2 SVHN results
Table 5: Test accuracy of SVHN on 10 clients using FedAvg with VGG16
@@ E.3 CIFAR-100 results
Table 6: Test accuracy of CIFAR-100 on 10 clients using FedAvg with ResNet-34
@@
Clean $0.0$ 74.15±0.11 71.58±0.31 68.31±1.97 72.08±1.8
Globalized Sym. $0.4$ 43.81±0.74 37.99±1.12 29.52 41.84
Asym. 44.5±0.6 40.48±0.44 36.61 41.47±1.56
Localized Sym. $\mathcal{U}(0.3,0.5)$ 44.53±0.9 36.28±3.51 28.52±1.32 42.01±3.32
Asym. 44.96±0.63 47.0±3.36 40.1±1.72 40.13±3.52
@@ F.1 CIFAR-10 results
Table 7: Test accuracy for globalized noise with different noise ratios on CIFAR-10 using FedAvg
@@
Table 8: Test accuracy for localized noise with different noise ratios on CIFAR-10 using FedAvg
@@
Sym. 0.1 $86.26\pm 0.08$ $53.31\pm 4.07$ $62.56\pm 11.59$
0.2 $80.91\pm 0.51$ $42.24\pm 1.59$ $47.57\pm 11.96$
0.3 $74.27\pm 0.70$ $31.02\pm 4.32$ $36.82\pm 10.65$
0.4 $66.05\pm 0.49$ $16.78\pm 1.58$ $27.21\pm 9.08$
0.5 $54.31\pm 1.96$ $13.39\pm 3.14$ $19.10\pm 7.17$
Asym. 0.1 $86.99\pm 0.18$ $59.34\pm 4.69$ $66.27\pm 9.20$
0.2 $80.32\pm 0.76$ $48.78\pm 1.81$ $58.49\pm 5.70$
0.3 $69.53\pm 0.39$ $36.21\pm 3.74$ $49.00\pm 8.07$
0.4 $56.96\pm 2.77$ $24.93\pm 3.18$ $39.80\pm 6.86$
0.5 $45.34\pm 1.06$ $23.09\pm 1.85$ $33.60\pm 6.62$
Table 9: Test accuracy for globalized noise with different noise ratios on MNIST using FedAvg
@@
Clean 0.0 99.17 ± 0.03 98.49 ± 0.29 98.05 ± 0.53
Sym. 0.1 95.67±0.07 74.95±1.03 75.4±5.73
0.2 89.63±0.13 57.84±0.93 58.88±5.83
0.3 81.2±0.02 46.72±1.55 49.09±6.4
0.4 71.49 39.88 41.30
0.5 60.03±0.6 33.57±1.45 34.62±4.48
0.6 48.97±1.74 29.5±0.84 27.68±2.78
0.7 37.33±0.9 22.37±0.88 22.86±2.78
Asym. 0.1 94.78±0.16 82.37±0.21 93.66±0.44
0.2 85.38±0.03 70.45±0.19 83.66±0.72
0.3 74.55±0.76 61.9±1.66 72.16±0.45
0.4 61.59±0.55 53.14±1.24 60.97±13.11
0.5 48.31±0.23 46.98±1.23 48.4±0.58
0.6 35.18±0.66 37.88±0.84 36.38±1.12
Table 10: Test accuracy for localized noise with different noise ratios on MNIST using FedAvg
@@
Sym. 0.1 95.26±0.54 82.79±1.95 80.89±3.16
0.2 89.36±0.73 67.62±2.32 62.81±4.95
0.3 80.96±1.14 54.45±1.75 50.75±4.91
0.4 70.99±0.43 42.5 40.75
0.5 59.69±1.42 32.19±1.23 32.21±3.91
Asym. 0.1 94.64±1.13 86.91±1.18 87.3±2.97
0.2 85.36±1.57 76.67±0.78 76.11±4.63
0.3 73.75±0.97 66.97±0.94 68.57±4.61
0.4 61.11±1.87 59.7±2.32 59.96
0.5 47.95±1.05 53.61±2.79 54.71±4.84
F.2 MNIST results
The partitions are iid, noniid-#label=3 and noniid-labeldir=0.1 partitions. Table 9 presents results under globalized noise with different noise ratios for 500 rounds of FedAvg training on 10 clients (the results for clean are using 250 rounds). Table 10 presents results under localized noise with different noise ratios for 500 rounds of FedAvg training on 10 clients.

### v4

@@ Appendix B Data Partitions
Given a data sample $(x,y)$, non-IID indicates the joint probability distribution $\mathcal{P}(x,y)$ is different for different client $i$ and $j$ ($i,j\in[K]$, $i\neq j$), that is, $\mathcal{P}_{i}(x,y)\neq\mathcal{P}_{j}(x,y)$. By Bayes’ theorem, $\mathcal{P}(x,y)$ can be rewritten as $\mathcal{P}(y|x)\mathcal{P}(x)$ and $\mathcal{P}(x|y)\mathcal{P}(y)$. Based on previous research (McMahan et al., 2017; Kairouz et al., 2021; Li et al., 2022a), we present IID and three representative non-IID schemes in FedNoisy. In this section, we first introduce IID partition, then partition methodologies for non-IID cases.
@@
IID   In IID, data distributions on each client are independent and identical. Following (McMahan et al., 2017), given a global dataset with $N$ training samples, we first randomly shuffle it, then partition it over $K$ clients, each receiving $\lfloor\frac{N}{K}\rfloor$ samples. Note that this partition keeps the original class ratio in the global dataset. If a class imbalance problem exists in the global dataset, then each client has the same class imbalance issue on its local dataset. We use iid to denote the IID partition.
@@
Quantity skew   In quantity distribution skew, different clients can hold different amounts of data, while keeping the same class distribution. For example, large academic hospitals usually have larger datasets compared with small community hospitals (Qu et al., 2021). Dirichlet distribution (Li et al., 2022a) and Log-Normal distribution (Acar et al., 2021) can be used to simulate this Non-IID partition. Here we use Dirichlet distribution following (Li et al., 2022a). Given $K$ clients, we draw $K$-dim vector $\mathbf{q}\sim\mathrm{Dir}_{K}(\alpha)$ and allocate $\lfloor q_{k}N\rfloor$ samples from global dataset $\mathcal{D}$ to client $k$ for $k=1,\ldots,K$. The parameter $\alpha>0$ can regulate the imbalance level of the quantity skew: a larger $\alpha$ will even lead to a partition result close to the IID partition. We use noniid-quantity=$\alpha$ to denote quantity distribution skew with parameter value $\alpha$.
@@
Label distribution skew   In label distribution skew, the prior distributions $\mathcal{P}(y)$ may vary across clients, while $\mathcal{P}(x|y)$ is shared in the federated system. For example, primary angle closure glaucoma (PACG) is much more common than primary open-angle glaucoma (POAG) in the Asian population than African, Hispanic population (Randhawa et al., 2021). In FedNoisy, we include two common practices for label distribution skew, namely distribution-based label distribution skew and quantity-based label distribution skew.
@@
  • Dirichlet label skew uses Dirichlet distribution to determine the class distribution on clients (Yurochkin et al., 2019; Wang et al., 2020b; Li et al., 2022a). Given $K$ clients and $C$ classes, we draw a $K$-dim vector $\mathbf{q}_{i}$ from $\mathrm{Dir}_{K}(\alpha)$ for each class $i\in[C]$ and allocate $q_{ik}$ proportion of global dataset with label $i$ to local client $k$. In addition to various overall sample sizes, the clients also have varying class distributions. A larger $\alpha$ encourages the class distribution among clients close to a uniform distribution. We use noniid-labeldir=$\alpha$ to denote distribution-based label distribution skew with parameter $\alpha$.
@@
  • Label-quantity skew assigns samples of specified number of classes on each client (McMahan et al., 2017; Li et al., 2022a; Nilsson et al., 2018). Here we follow the strategy in (Li et al., 2022a). Given the number of classes on each client as $c$, we first assign $c$ different classes to each client and then equally divide samples of each class according to the number of clients assigned to them. When $c$ equals the total class number in the federation, the partition is equivalent to the IID partition. We use noniid-#label=$c$ to denote quantity-based label distribution skew with parameter $c$.
@@ Appendix C Fair comparison between globalized and localized noise scene
Figure 8. Overall noisy label ratio in localized symmetric noise $\mathcal{U}(\varepsilon_{local}-\sigma,\varepsilon_{local}+\sigma)$ on CIFAR-10 with 10 clients. From left to right: iid, noniid-labeldir, noniid-#label and noniid-labeldir partition.
@@ Appendix D Gradient norm under localized noise setting
Figure 9. Gradient norm of global model for clean and localized noise $\varepsilon_{k}\sim\mathcal{U}(0.3,0.5)$.
@@ Appendix E Results on other datasets
E.1. MNIST results on FedAvg
@@
Table 5. Test accuracy of MNIST on 10 clients using FedAvg with CNN
@@ E.2. SVHN results
Table 6. Test accuracy of SVHN on 10 clients using FedAvg with VGG16
@@ E.3. CIFAR-100 results
Table 7. Test accuracy of CIFAR-100 on 10 clients using FedAvg with ResNet-34
@@
Clean $0.0$ $74.15\pm 0.11$ $71.58\pm 0.31$ $68.31\pm 1.97$ $72.08\pm 1.80$
Globalized Sym. $0.4$ $43.81\pm 0.74$ $37.99\pm 1.12$ $29.2\pm 0.64$ $42.51\pm 2.68$
Asym. $44.50\pm 0.60$ $40.48\pm 0.44$ $36.35\pm 0.48$ $41.47\pm 1.56$
Localized Sym. $\mathcal{U}(0.3,0.5)$ $44.53\pm 0.9$ $36.28\pm 3.51$ $28.52\pm 1.32$ $42.01\pm 3.32$
Asym. $44.96\pm 0.63$ $47.00\pm 3.36$ $40.10\pm 1.72$ $40.13\pm 3.52$
E.4. Clothing1M results
FedAvg results on Clothing1M using ResNet-50 (He et al., 2016), 150 rounds, 5 local epochs, SGD, 10 clients with $100\%$ selection ratio. The partition setting is iid, noniid-#label=5, noniid-labeldir=0.1, noniid-quantity=0.1. Results are shown in Table 8.
Table 8. Test accuracy of Clothing1M on 10 clients using FedAvg with ResNet-50
Noise mode Data partition
iid noniid-#label noniid-labeldir noniid-quantity
Real $70.99\pm 0.33$ $53.68\pm 5.26$ $48.41\pm 9.52$ $67.70\pm 0.87$
@@ F.1. CIFAR-10 results
Table 9. Test accuracy for globalized noise with different noise ratios on CIFAR-10 using FedAvg
@@
Table 10. Test accuracy for localized noise with different noise ratios on CIFAR-10 using FedAvg
@@
Sym. $\mathcal{U}(0.0,0.2)$ $86.26\pm 0.08$ $53.31\pm 4.07$ $62.56\pm 11.59$
$\mathcal{U}(0.1,0.3)$ $80.91\pm 0.51$ $42.24\pm 1.59$ $47.57\pm 11.96$
$\mathcal{U}(0.2,0.4)$ $74.27\pm 0.70$ $31.02\pm 4.32$ $36.82\pm 10.65$
$\mathcal{U}(0.3,0.5)$ $66.05\pm 0.49$ $16.78\pm 1.58$ $27.21\pm 9.08$
$\mathcal{U}(0.4,0.6)$ $54.31\pm 1.96$ $13.39\pm 3.14$ $19.10\pm 7.17$
Asym. $\mathcal{U}(0.0,0.2)$ $86.99\pm 0.18$ $59.34\pm 4.69$ $66.27\pm 9.20$
$\mathcal{U}(0.1,0.3)$ $80.32\pm 0.76$ $48.78\pm 1.81$ $58.49\pm 5.70$
$\mathcal{U}(0.2,0.4)$ $69.53\pm 0.39$ $36.21\pm 3.74$ $49.00\pm 8.07$
$\mathcal{U}(0.3,0.5)$ $56.96\pm 2.77$ $24.93\pm 3.18$ $39.80\pm 6.86$
$\mathcal{U}(0.4,0.6)$ $45.34\pm 1.06$ $23.09\pm 1.85$ $33.60\pm 6.62$
Table 11. Test accuracy for globalized noise with different noise ratios on MNIST using FedAvg
@@
Clean 0.0 $99.17\pm 0.03$ $98.49\pm 0.29$ $98.05\pm 0.53$
Sym. 0.1 $95.67\pm 0.07$ $74.95\pm 1.03$ $75.4\pm 5.73$
0.2 $89.63\pm 0.13$ $57.84\pm 0.93$ $58.88\pm 5.83$
0.3 $81.2\pm 0.02$ $46.72\pm 1.55$ $49.09\pm 6.4$
0.4 $71.75\pm 0.67$ $39.02\pm 1.76$ $40.49\pm 5.57$
0.5 $60.03\pm 0.6$ $33.57\pm 1.45$ $34.62\pm 4.48$
0.6 $48.97\pm 1.74$ $29.5\pm 0.84$ $27.68\pm 2.78$
0.7 $37.33\pm 0.9$ $22.37\pm 0.88$ $22.86\pm 2.78$
Asym. 0.1 $94.78\pm 0.16$ $82.37\pm 0.21$ $93.66\pm 0.44$
0.2 $85.38\pm 0.03$ $70.45\pm 0.19$ $83.66\pm 0.72$
0.3 $74.55\pm 0.76$ $61.9\pm 1.66$ $72.16\pm 0.45$
0.4 $61.59\pm 0.55$ $53.14\pm 1.24$ $60.97\pm 13.11$
0.5 $48.31\pm 0.23$ $46.98\pm 1.23$ $48.4\pm 0.58$
0.6 $35.18\pm 0.66$ $37.88\pm 0.84$ $36.38\pm 1.12$
Table 12. Test accuracy for localized noise with different noise ratios on MNIST using FedAvg
@@
Sym. $\mathcal{U}(0.0,0.2)$ $95.26\pm 0.54$ $82.79\pm 1.95$ $80.89\pm 3.16$
$\mathcal{U}(0.1,0.3)$ $89.36\pm 0.73$ $67.62\pm 2.32$ $62.81\pm 4.95$
$\mathcal{U}(0.2,0.4)$ $80.96\pm 1.14$ $54.45\pm 1.75$ $50.75\pm 4.91$
$\mathcal{U}(0.3,0.5)$ $70.99\pm 0.43$ $42.73\pm 0.96$ $40.64\pm 3.93$
$\mathcal{U}(0.4,0.6)$ $59.69\pm 1.42$ $32.19\pm 1.23$ $32.21\pm 3.91$
Asym. $\mathcal{U}(0.0,0.2)$ $94.64\pm 1.13$ $86.91\pm 1.18$ $87.30\pm 2.97$
$\mathcal{U}(0.1,0.3)$ $85.36\pm 1.57$ $76.67\pm 0.78$ $76.11\pm 4.63$
$\mathcal{U}(0.2,0.4)$ $73.75\pm 0.97$ $66.97\pm 0.94$ $68.57\pm 4.61$
$\mathcal{U}(0.3,0.5)$ $61.11\pm 1.87$ $59.70\pm 2.32$ $61.06\pm 4.21$
$\mathcal{U}(0.4,0.6)$ $47.95\pm 1.05$ $53.61\pm 2.79$ $54.71\pm 4.84$
F.2. MNIST results
The partitions are iid, noniid-#label=3 and noniid-labeldir=0.1 partitions. Table 11 presents results under globalized noise with different noise ratios for 500 rounds of FedAvg training on 10 clients. Table 12 presents results under localized noise with different noise ratios for 500 rounds of FedAvg training on 10 clients. Figure 10 presents the plot of the accuracy drop ratio between Non-IID and IID with different noise ratios.
Figure 10. Accuracy drop ratio from IID to Non-IID with globalized noise on MNIST. Give noise ratio $\varepsilon$, the accuracy drop ratio of certain non-iid partition at $\varepsilon$ is calculated by $[\mathrm{acc}_{\texttt{iid}}(\varepsilon)-\mathrm{acc}_{\texttt{non-iid}}(\varepsilon)]/\mathrm{acc}_{\texttt{iid}}(\varepsilon)$. The dotted lines mark the values when $\varepsilon=0.0$, and the dash-dotted lines mark accuracy drop ratio equals 0. $\varepsilon$ is $\varepsilon_{global}$ for globalized noise, and $\varepsilon_{local}$ for localized noise $\mathcal{U}(\varepsilon_{local}-0.1,\varepsilon_{local}+0.1)$. From left to right: globalized noise, localized noise.
Appendix G More Baseline Results
The visualization of accuracy and computation trade-off on MNIST is shown in Figure 11.
Figure 11. The trade-off between accuracy and per-round local computation time on MNIST on 10 clients under noniid-#label=3 partition with symmetric noise. From left to right: globalized noise, localized noise.